It has been quiet during the day at the HET this week with day staff on holiday. If the weather clears the night staff will be plenty busy working through the queue. The current weather outlook for the week is cloudy with possible showers. Monday night we were able to do science through thin clouds and successfully collected data for many programs that do not require perfect seeing or zero cloud cover. The bulk of the science data collected in the beginning of the week has been with LRS2-B and LRS2-R. Fingers crossed the weather clears as we approach the new moon on Thursday.